On the first of three tests, Charlotte scored 72 points. On the third test, her score was one point more than on her second test. Her average on the three tests were 83. What were her scores on the second and third tests?
Answer by mananth(16946)

On the first of three tests, Charlotte scored 72 points. On the third test, her score was one point more than on her second test. Her average on the three tests were 83. What were her scores on the second and third tests?
..
average on 3 tests = 83
so total = 249
..
first = 72
second x
third = x+1
...
x+x+1+72=249
2x+73=249
2x=249-73
2x=176
/2
x=88 second test
third = 89
...
m.ananth@hotmail.ca